OpenStudy (anonymous):

Fill in the missing term so that the quadratic equation has a graph that opens up, with a vertex of (– 2, – 16), and x intercepts at x = -6 and x = 2. (Do not include the negative sign in your answer.) y = x2 + 4x - ___

OpenStudy (campbell_st):

if it has intercepts at x = -6 and x = 2 and a vertex at (-2, -16) the parabola is concave up. then the equation is y= (x +6)(x -2) expanding gives y = x^2 + 4x - 12

OpenStudy (campbell_st):

the missing value is 12